OpenStudy (anonymous):

Which of the following is a linear function? y = x3 + 1 5/6 x = y – 4 y = 8x^2 y = 1/5x

OpenStudy (anonymous):

Should this question have multiple answers? Because both b and d are linear with x. Meaning, if you were to graph them, they would be lines! a is a cubic, and is not a line. c is a parabola!

Directrix (directrix):

Grouping symbols are missing which lead to confusion. Something like: 5/6 x = y – 4 Is that (5/6) * x = y - 4 OR is that 5 / (6x) = y - 4 @Samirah16 Click on the blue DRAW tab and sketch by hand the equations as they appear in your text. That will clear the confusion. Thanks.

OpenStudy (anonymous):